# Cocoa: Co-Planning and Co-Execution with AI Agents

### v2

@@ Abstract.
We present Cocoa, a system that implements a novel interaction design pattern—interactive plans—for users to collaborate with an AI agent on complex, multi-step tasks in a document editor. Cocoa harmonizes human and AI efforts and enables flexible delegation of agency through two actions: Co-planning (where users collaboratively compose a plan of action with the Agent) and Co-execution (where users collaboratively execute plan steps with the Agent). Using scientific research as a sample domain, we motivate the design of Cocoa through a formative study with 9 researchers while also drawing inspiration from the design of computational notebooks. We evaluate Cocoa through a user study with 16 researchers and find that when compared to a strong chat baseline, Cocoa improved agent steerability without sacrificing ease of use. A deeper investigation of the general utility of both systems uncovered insights into usage contexts where interactive plans may be more appropriate than chat, and vice versa. Our work surfaces numerous practical implications and paves new paths for interactive interfaces that foster more effective collaboration between humans and agentic AI systems.
copyright: acmcopyrightjournalyear: 2025doi: xxxxxxx.xxxxxxxconference: arXiv; January; 2025price: 15.00isbn: xx-x-xxx-xxxx-xxxxx
@@ 1. Introduction
Figure 2. An overview of the Cocoa user interface. An interactive plan (A) affords human-AI co-planning and co-execution: a researcher and the AI agent can collaboratively edit the plan in the document and execute the plan steps, similar to executing code cells in a computational notebook. Steps can be assigned to the AI agent (B) or the researcher (C). The researcher can freely edit the AI agent’s outputs in an interactive sidebar (D) and add new information (E) to help steer the agent with their feedback and expertise. In this example, the first three steps of a plan to summarize methods for human-feedback elicitation have already been executed, and the agent is requesting guidance from the user in the next step.
Since the advent of personal computing, researchers and practitioners in artificial intelligence (AI) and human-computer interaction (HCI) have set sights on developing intelligent AI agents that can help perform everyday computer-based tasks on our behalf (Maes, 1994; Maes and Wexelblat, 1996; Lieberman, 1997; Horvitz, 1999). Thanks to recent advancements in large language models (LLMs) and LLM agent frameworks involving reasoning, planning, memory, and tool use (Yao et al., 2022b; Sumers et al., 2023; Schick et al., 2023; Wang et al., 2023; Kim et al., 2024), AI agents can now shop online (Yao et al., 2022a), write software (Yang et al., 2024; Jimenez et al., 2023; Wu, 2024), participate in auctions (Chen et al., 2023), and perform other multi-step, computer-based tasks (Anthropic, 2024). These developments demand deeper explorations into strategies for fostering synergistic collaboration between human users and AI agents (Kapoor et al., 2024; Collins et al., 2024; Ibrahim et al., 2024). After all, effective interaction with agents [^1] Henceforth, our use of the term “agents” will refer to AI agents rather than human agents. in the real world will need to make use of techniques that use human guidance to significantly improve agent performance and utility, and better align with users’ nuanced needs and preferences (Kapoor et al., 2024; Shi et al., 2024). However, explorations of such techniques are still nascent. Much of modern agent development efforts focus on autonomous agents without the need for human supervision, through techniques such as chain-of-thought reasoning (Wei et al., 2022) and interleaving reasoning traces and actions (Yao et al., 2022b).
@@
While creating fully autonomous agents may be compelling in theory, it has notable limitations in practice. First, we cannot steer the agent with our expertise and worldly understanding. Two salient opportunities for agent steering in task completion workflows which we focus on in this work are planning and execution. Agents’ abilities to generate executable plans of action have been shown to be unreliable across many domains (Valmeekam et al., 2023). Without a quality plan, agents can easily veer off track, wasting resources without achieving meaningful results. When executing the plan, human guidance can significantly boost agent performance—Shi et al. (Shi et al., 2024) found that even simple feedback from programmers to an agent for solving Olympiad-style programming problems increased agent accuracy from 0% to over 85%. Second, and perhaps more importantly, removing human agency in favor of AI agency can pose heightened safety risks, disempower us to think critically and be creative, and harm our well-being more generally (Chan et al., 2023; Bennett et al., 2023; Chan et al., 2024; Collins et al., 2024). For many nuanced, subjective, and high-stakes tasks, human input must be considered for AI-powered systems to be successful and aligned with users’ personal needs and goals (Cheong et al., 2024; Feng et al., 2024; Terry et al., 2023).
When using LLMs through chat interfaces, a user can engage in dynamic task planning and execution by sending, receiving, and evaluating content generated by the model. However, prior work in HCI and human-AI interaction has suggested that planning-enriched workflows—where the LLM first decomposes a high-level task into lower-level ones and shows them to the user before taking action—can improve usability and transparency (Ma et al., 2023; Lawley and MacLellan, 2024). In contrast, rather than automatically generating plans, researchers have also developed systems for humans to manually compose plans and then use those plans to guide LLM generation (Yen et al., 2024). However, chat interfaces may not be well-suited for these workflows—consequently, emerging efforts have started to explore alternative interaction paradigms when working with agents to perform complex, multi-step, and often long-running tasks (Ma et al., 2023; Suh et al., 2023; Pu et al., 2024; OpenAI, 2024). For example, researchers have engineered node-based canvas interfaces for information sensemaking (Suh et al., 2023), ideation (Pu et al., 2024), and even task decomposition (Wu et al., 2021) with LLMs. While canvases provide freedom for unstructured exploration with AI models, their lack of structure can hinder effective orchestration of efforts and agency between humans and AI when working through structured, sequential plans of action. We see opportunities for new interaction design patterns to serve as shared operational artifacts to bridge human and AI planning, such that it facilitates collaborative human-AI plan composition and execution.
@@
In this work, we introduce Cocoa, a novel system for Co-planning and Co-execution with AI Agents. Cocoa embeds AI gents into a document editor—a common site for planning—through a new interaction design pattern which we call interactive plans. Interactive plans orchestrate actions between a human user and an AI agent and enable flexible delegation of human and AI agency. As a result, Cocoa supports human-AI co-planning: the agent, when invoked in the document, will first propose a plan of action that the user can freely modify. This plan is interactive and seamlessly integrated into the document—the user can then edit the plan steps through familiar interactions that mirror typical document editing and assign steps to the agent or themselves. Cocoa also supports human-AI co-execution: drawing inspiration from the design of computational notebooks, the interactive plan allows the user and the agent to collaboratively complete one step at a time or re-execute steps as desired. The user can interactively refine the agent’s intermediate outputs and also manually take over steps themselves to steer the workflow in a more effective direction. Most importantly, Cocoa interleaves co-planning and co-execution, such that users can smoothly transition between the two and modify their plans based on outputs from execution, and vice versa.
We motivate the design and development of Cocoa by focusing on scientific research as a use case, due to the complex, multi-step, and information-dense workflows researchers often engage in throughout their work (Kang et al., 2022b; Fok et al., 2022, 2023; Kang et al., 2022c; Lee et al., 2024b). Through formative studies with 9 researchers, we learned that their real-world research project documents—a running document where researchers informally jot down ideas, open questions, meeting notes, and more—contain rich externalizations of researchers’ (often preliminary) reasoning and thus serve as ideal environments for interacting with an agent to further tackle research ideas. We also gathered examples of participants’ plans and preferences for tasks they would delegate to AI to inform the behavior of Cocoa. After developing Cocoa, we evaluated it against a strong chat baseline—a more familiar interface powered by the same agent—through a user study with 16 researchers. We found that compared to our baseline, Cocoa enabled users to better steer the agent in more helpful directions without sacrificing ease of use. A deeper investigation into qualitative data from the study revealed that the general utility of both systems depended on the nature of the task, and we unpack scenarios in which using interactive plans as a design pattern may be more appropriate than chat interfaces, and vice versa. We conclude with a discussion of how agent interfaces may simultaneously capture benefits of interactive plans and chat, practical reasons—namely cost and safety—for including user steps within interactive plans, and the broader applicability of interactive plans outside of the document context.
@@
    A formative study with 9 researchers that uncovered user needs and opportunities to better support planning and execution in research project documents.
@@
    Qualitative data from the above provide valuable insights into when interactive plans may be preferred over chat interfaces, and vice versa.
  5. (5)
    An in-depth discussion of our work’s practical implications, including how interactive plans may be used in diverse settings beyond documents, and the trade-offs when compared with chat-based interfaces.
@@ 2.1. Planning and Interactivity in LLM Agents
Prior work has envisioned autonomous software agents that can assist users in various ways, from performing complex tasks on the user’s behalf to monitoring events and procedures (Maes, 1994). These “interface agents” (Maes and Wexelblat, 1996; Lieberman, 1997) can perform visible actions in a direct manipulation interface without explicit instructions from the user. Two decades later, the same vision endures with LLM agents. Core components in LLM agent architectures typically consist of memory, reasoning, planning, and tool use (Yao et al., 2022b; Zhou et al., 2023; Liu et al., 2024b; Sumers et al., 2023; Wang et al., 2023; Kim et al., 2024).
@@
Central to LLM agents’ operations is multi-step reasoning, often implemented by chain-of-thought (CoT) (Wei et al., 2022; Qin et al., 2023; Yao et al., 2022b) and subsequent tree-based methods (Yao et al., 2023; Zhuang et al., 2023). CoT provides a series of intermediate reasoning steps as exemplars in prompting to boost performance on complex reasoning tasks (Wei et al., 2022). CoT [^2] Henceforth, we use CoT as an umbrella term encompassing chain-of-thought, trees-of-thought (Yao et al., 2023), and related methods. is crucial for facilitating task decomposition and therefore LLM agents’ planning capabilities (Qin et al., 2023; Chen et al., 2023). However, even with CoT, critical investigations of LLMs’ abilities to autonomously generate executable plans reveal an average success rate of only 12% across diverse domains (Valmeekam et al., 2023). Indeed, although CoT can help improve model planning for tasks with a well-defined solution and logical steps to get to that solution—e.g., numerical, tabular, and knowledge-based reasoning—(Kim et al., 2024), it may not be so effective in domains that require high-expertise, tacit knowledge to navigate towards ambiguous solutions (Hao et al., 2023). Unlike model properties that show empirical improvement through scaling laws, limitations of planning in these domains may not resolve with scale alone as 1) tacit knowledge is not well-documented in training data and is thus difficult for a model to grasp (Cheong et al., 2024; Feng et al., 2023; Zhang et al., 2024), and 2) there may be no “correct” workflow for CoT to follow and verify the correctness post-hoc (Jirotka et al., 2013). Scientific research is one such domain (Zhang et al., 2024).
@@
In light of this, recent work has recognized the need to interactively incorporate user feedback for improving LLM agents’ planning capabilities and beyond (Valmeekam et al., 2023; Yao et al., 2022b; Kazemitabaar et al., 2024; Lawley and MacLellan, 2024; Shao et al., 2024), particularly within scientific discovery (Majumder et al., 2024). For example, Lawley and MacLellan (Lawley and MacLellan, 2024) architected an approach where user interaction is used to guide the model in planning for unseen tasks on-the-fly using a hierarchical network of smaller actions. Yet, developments in this area have been scant, despite the new challenges identified for human-agent communication (Bansal et al., 2024). Interfaces for LLM assistants (e.g., AgentGPT (Reworkd, 2024), OpenAI Assistants Playground (OpenAI Platform, 2024), Devin (Wu, 2024)) have been primarily limited to text-heavy chat interfaces targeted at monitoring agent activity rather than empowering the user to proactively collaborate with the agent. A small handful of works have started to experiment with interactive techniques for LLM long-running agents. For instance, Kazemitabaar et al. (Kazemitabaar et al., 2024) developed interfaces for data analysts to edit an execution plan of an LLM to provide more control points for steering behavior, while NL2Plan (Gestrin et al., 2024) accepts a bulleted list of user suggestions for the agent as feedback. However, these interactions are corrective rather than collaborative—that is, users would typically engage in these interactions to correct agent behavior post-hoc rather than proactively iterate back-and-forth with the agent at multiple points in the workflow. Rather than treating human input as a corrective mechanism, we aim to center human agency through the latter.
The GUI revolutionized computing by shifting interactions from text-based command lines to direct manipulation of representations aligned with our mental models (Hutchins et al., 1985; Horvitz, 1999). We see a need for a similar transformation in LLM agents—from autonomous agents to interactive agents. By “interactive agents,” we mean AI agents that can effectively elicit and incorporate key guidance from the user through interactive artifacts shared and co-created with users. For scientific co-planning, we are particularly interested in embedding interactive agents into a document environment, as such environments are common for planning and organizing research projects while offering valuable context for an agent.
For decades, HCI researchers have been exploring the interplay between human agency—the feeling of control over one’s actions and ability to act in accordance with personal goals and values (Bennett et al., 2023)—and machine agency—the ability for a technical system to complete a task, often proactively, with minimal human supervision (Chan et al., 2023). Horvitz (Horvitz, 1999) described this interplay as mixed-initiative, prompting the establishment of a wide range of guidelines for more effective human-AI interaction (Amershi et al., 2019; Microsoft, 2024; Google, 2021; Apple, 2023). The field of embodied AI also explores this through the notion of “shared autonomy” (Fontaine and Nikolaidis, 2022; Nikolaidis et al., 2017). Our work introduces an interaction design pattern for LLM agents that offers a shared, collaborative artifact that aligns agent workflows and user mental models to facilitate effective collaboration between users and AI agents.
@@ 2.2. Computational Notebooks
The computational notebook is an interactive paradigm that organizes code, data, rich outputs (e.g., tables, visualizations, interactive widgets), and other artifacts into linearly arranged cells (Chattopadhyay et al., 2020; Lau et al., 2020), reflecting Donald Knuth’s early visions of literate programming (Knuth, 1984). Computational notebooks are a well-studied paradigm in HCI (e.g., (Choi et al., 2023; Ayobi et al., 2023; Zheng et al., 2022; Lin et al., 2023; Chattopadhyay et al., 2020; O’Leary et al., 2023)). Lau et al. (Lau et al., 2020) analyzed 60 notebook systems in academia and industry to formulate a design space for computational notebooks based on dimensions including data sources, editor style, and execution order. Informed by these prior work, of particular interest to us are design considerations that impact notebooks’ interactive properties and general utility.
@@
In our work, we take inspiration from design decisions made for computational notebooks because of the many parallels between workflows in notebooks and mixed-initiative task completion systems (see Section LABEL:s:rw-task-completion). For instance, when data scientists often break down a high-level data analysis task into executable cells in computational notebooks (Kery et al., 2018; Rule et al., 2018). As they execute each cell and inspect outputs, they can verify the quality and sensibility of outputs before moving forward in their analysis. By drawing these parallels, we can reveal new opportunities to design interactive workflows and interfaces for LLM agents. Just like how a data scientist may easily add new cells in a notebook or reconfigure existing cells to adapt their analysis plan, a user may interactively edit an agent’s plan of execution to better steer the agent in productive directions. Furthermore, an agentic system may simultaneously be more usable and resource-efficient if a user could direct an agent to iterate on a subtask to improve its output before continuing onto subtasks dependent on that output. This also demands new ways of viewing and editing agent outputs on subtasks, which increases opportunities for human input over existing approaches such as simply logging agent actions (AutoGPT, 2024; Reworkd, 2024; OpenAI Platform, 2024) Our work exploits these parallels to contribute new design patterns for steering agentic AI systems.
2.3. AI-Powered Tools for Scientific Research
Significant efforts in recent years have advanced how AI can augment scientific research. These efforts include tools for paper reading (Chang et al., 2023; Rachatasumrit et al., 2022) and skimming (Fok et al., 2022), literature review (Lee et al., 2024b; Palani et al., 2023), paper recommendation (Kang et al., 2022b; Cohan et al., 2020; Kang et al., 2022a; Soliman et al., 2024), information retrieval and sensemaking (Chan et al., 2018; Kang et al., 2023; Fok et al., 2023; Wang et al., 2024), as well as toolkits (Lo et al., 2023) and academic search engines (for AI, 2024; Google, 2024; Elicit, 2024) that power or combine these tools. More generally, progress in this area has given rise to excitement for a “computational inflection for scientific discovery” (Hope et al., 2022).
Of particular interest to us are works that leverage these recent advances to help researchers with literature-augmented planning. Planning a research project is a cognitively demanding, complex process consisting of iterative cycles of divergent and convergent thinking grounded in literature review (Palani et al., 2023; Dean et al., 2006; Chang et al., 2023). Research prototypes have been developed to indirectly assist with planning in two primary ways. Some prototypes help researchers interactively make sense of papers to orient their work (Chang et al., 2023; Lee et al., 2024b; Palani et al., 2023; Kang et al., 2023; Fok et al., 2023). Other prototypes have also attempted to generate high-fidelity research ideas directly (Gu and Krenn, 2024; Baek et al., 2024; Jansen et al., 2024). However, when these ideas were evaluated by human researchers, they received dismal scores. On a 5-point Likert scale (1 = not interesting, 5 = very interesting), Gu et al. found that 75% of generated suggestions from their study received a rating of 3 or lower, with 37% receiving a rating of 1 (Gu and Krenn, 2024). Similarly, Baek et al. saw limited enthusiasm from expert human annotators for AI-generated ideas from their agent (Baek et al., 2024).

### v3

@@ Abstract.
Human collaboration benefits from continuous coordination— planning, delegating tasks, sharing progress, and adjusting objectives—to align on shared goals. However, agentic AI systems often limit users to previewing or reviewing an agent’s plans for fully autonomous execution. While this may be useful for confirmation and correction, it does not support deeper collaboration between humans and AI agents. We present Cocoa, a system that introduces a novel design pattern—interactive plans—for collaborating with an AI agent on complex, multi-step tasks. Informed by a formative study ($n=9$), Cocoa builds on interaction designs from computational notebooks and document editors to support flexible delegation of agency through Co-planning and Co-execution, where users collaboratively compose and execute plans with an Agent. Using scientific research as a sample domain, our lab ($n=16$) and field deployment ($n=7$) studies found that Cocoa improved agent steerability without sacrificing ease-of-use compared to a strong chat baseline. Additionally, researchers valued Cocoa for real-world projects and saw the interleaving of co-planning and co-execution as an effective novel paradigm for human-AI collaboration.
copyright: acmcopyrightjournalyear: 2025doi: xxxxxxx.xxxxxxxconference: arXiv; April; 2025price: 15.00isbn: xx-x-xxx-xxxx-xxxxx
@@ 1. Introduction
Figure 2. An overview of the Cocoa user interface. An interactive plan (A) affords human-AI co-planning and co-execution: a researcher and the AI agent can collaboratively edit the plan in the document and execute the plan steps, similar to executing code cells in a computational notebook. Steps can be assigned to the AI agent (B) or the researcher (C). The researcher can freely edit the AI agent’s outputs in an interactive sidebar (D), such as adding relevant papers that the agent did not find (E) to help steer the agent with their feedback and expertise. In this example, the first three steps of a plan to summarize methods for human feedback elicitation have already been executed, and the agent is requesting guidance from the user in the next step.
[…]
An overview of the Cocoa user interface, which resembles a document editor. An interactive plan affords human-AI co-planning and co-execution: a researcher and the AI agent can collaboratively edit the plan in the document and execute the plan steps, similar to executing code cells in a computational notebook. Steps can be assigned to the AI agent or the researcher. The researcher can freely edit the AI agent’s outputs in an interactive sidebar, such as adding relevant papers that the agent did not find, to help steer the agent with their feedback and expertise. In the example shown, the first three steps of a plan to summarize methods for human feedback elicitation have already been executed, and the agent is requesting guidance from the user in the next step.
Since the advent of personal computing, researchers and practitioners in artificial intelligence (AI) and human-computer interaction (HCI) have set sights on developing intelligent AI agents that can help perform everyday computer-based tasks on our behalf (Maes, 1994; Maes and Wexelblat, 1996; Lieberman, 1997; Horvitz, 1999). Recent advancements in large language models (LLMs) and agent frameworks involving reasoning, planning, memory, and tool use (Yao et al., 2022b; Sumers et al., 2023; Schick et al., 2023; Wang et al., 2023; Kim et al., 2024) have accelerated the development of fully autonomous AI agents that can conduct complex tasks with varying degrees of success. These tasks include shopping online (Yao et al., 2022a), writing software (Yang et al., 2024; Jimenez et al., 2023; Wu, 2024), performing “deep research” to assemble a report based on web sources (Google, 2025; OpenAI, 2025a; Perplexity, 2025), and general computer-based tasks (Anthropic, 2024; OpenAI, 2025b). By contrast, turn-based and direct manipulation LLM systems that shift the burden of task planning and progress monitoring to users—such as chat or node-based interfaces (Suh et al., 2023; Liu et al., 2024a; Wu et al., 2022; Pu et al., 2024)—have become popular in HCI in recent years for offering users greater agency and control.
While some systems have attempted to balance AI and human agency by allowing users to review an agent’s reasoning chains after execution for verification (OpenAI, 2024; Guo et al., 2025), or to edit and confirm its plans before execution (Google, 2025; OpenAI, 2025b), few works have explored how to deeply involve and collaborate with users during iterative planning and execution. We argue that effective human interaction with agents remains essential for agents to operate effectively and safely in the real world (Kapoor et al., 2024; Shi et al., 2024), such that human-AI teams can accomplish tasks that cannot easily be completed by humans or AI alone (Bansal et al., 2021). Realizing this vision demand deeper explorations into strategies for fostering synergistic collaboration between human users and AI agents (Kapoor et al., 2024; Collins et al., 2024; Ibrahim et al., 2024).
@@
More concretely, this lack of human-AI collaboration in task planning and execution can lead to several issues. First, we cannot steer the agent with our expertise and worldly understanding. In domains where agents can not yet produce a quality plan and the ability to steer agents towards one, agents can easily veer off track, wasting resources without achieving meaningful results (Valmeekam et al., 2023; Husain et al., 2025). When executing the plan, human guidance can significantly boost agent performance—Shi et al. (Shi et al., 2024) found that even simple feedback from programmers to an agent for solving Olympiad-style programming problems increased agent accuracy from 0% to over 85%. Second, and perhaps more importantly, removing human agency in favor of AI agency can pose heightened safety risks, disempower us to think critically and be creative, and harm our well-being more generally (Chan et al., 2023; Bennett et al., 2023; Kulveit et al., 2025; Collins et al., 2024). More fundamentally, for many nuanced, subjective, and high-stakes tasks, human input must be considered for AI agents to be successful and aligned with users’ personal needs and goals (Cheong et al., 2024; Feng et al., 2024; Terry et al., 2023).
@@
In this work, we introduce Cocoa, a novel system for Co-planning and Co-execution with AI Agents. We chose scientific research tasks as a sample domain due to the complex, multi-step, and information-dense workflows researchers often engage in throughout their work (Kang et al., 2022b; Fok et al., 2022, 2023; Kang et al., 2022c; Lee et al., 2024b). Cocoa embeds AI agents into a document editor—a common environment many researchers use to keep track of tasks and progress for their research projects. Taking inspiration from computational notebook interfaces where data scientists compose program instructions and manage their execution and outputs, we introduce a new interaction design pattern for AI agents called interactive plans. Interactive plans orchestrate actions between a human user and an AI agent and enable flexible delegation of human and AI agency. As a result, Cocoa supports human-AI co-planning: the agent, when invoked in the document, will first propose a plan of action that is interactive and seamlessly integrates into the document. The user can then edit the plan steps through familiar interactions that mirror typical document editing and assign steps to the agent or themselves. Cocoa also supports human-AI co-execution: drawing inspiration from the design of computational notebooks, the interactive plan allows the user and the agent to collaboratively complete one step at a time or re-execute steps as desired. The user can interactively refine the agent’s intermediate outputs and also manually take over steps themselves to steer the workflow in a more effective direction. Most importantly, Cocoa interleaves co-planning and co-execution, such that users can smoothly transition between the two and modify their plans based on outputs from execution, and vice versa.
@@
    A formative study with 9 researchers with a focus around their real-world project documents that uncovered needs and opportunities to better support planning and execution.
@@
    A 7-day deployment study of Cocoa with 7 researchers that offers insights into how interactive plans can provide value to researchers in their day-to-day work.
@@ 2.1. Planning and Interactivity in LLM Agents
Prior work has identified core components in LLM agent architectures to consist of memory, reasoning, planning, and tool use (Yao et al., 2022b; Zhou et al., 2023; Liu et al., 2024b; Sumers et al., 2023; Wang et al., 2023; Kim et al., 2024). Central to LLM agents’ longitudinal operations is long-term planning, which demands capable reasoning abilities (Hao et al., 2023) and thus is often implemented by chain-of-thought (CoT) [^1] We use CoT as an umbrella term encompassing chain-of-thought, trees-of-thought (Yao et al., 2023; Zhuang et al., 2023), and related methods. (Wei et al., 2022; Qin et al., 2023; Yao et al., 2022b; OpenAI, 2024). CoT provides a series of intermediate reasoning steps as exemplars in prompting to boost performance on complex reasoning tasks (Wei et al., 2022; OpenAI, 2024). CoT is crucial for facilitating task decomposition and therefore LLM agents’ planning capabilities (Qin et al., 2023; Chen et al., 2023). However, even with CoT, critical investigations of LLMs’ abilities to autonomously generate executable plans reveal limited success across diverse domains (Valmeekam et al., 2023). Indeed, although CoT can help improve model planning for tasks with well-defined, objective solutions that provide clear signals for reinforcement learning—e.g., numerical, tabular, and knowledge-based reasoning (Kim et al., 2024)—it may not be so effective in domains that require expert tacit knowledge to navigate ambiguous problem spaces with no single right answer (Hao et al., 2023). Unlike model properties that show empirical improvement through scaling laws, limitations of planning in these domains may not resolve with scale alone as 1) tacit knowledge is not well-documented in training data and is thus difficult for a model to robustly learn (Cheong et al., 2024; Feng et al., 2023; Zhang et al., 2024), and 2) there may be no “correct” workflow for CoT to follow and verify the correctness post-hoc (Jirotka et al., 2013). Scientific research is one such domain (Zhang et al., 2024).
@@
In light of this, recent work has recognized the need to interactively incorporate user feedback for improving LLM agents’ planning capabilities and beyond (Valmeekam et al., 2023; Yao et al., 2022b; Kazemitabaar et al., 2024; Lawley and MacLellan, 2024; Shao et al., 2024; OpenAI, 2025b), particularly within scientific discovery (Majumder et al., 2024; Gottweis et al., 2025). For example, Lawley and MacLellan (Lawley and MacLellan, 2024) architected an approach where user interaction is used to guide the model in planning for unseen tasks on-the-fly using a hierarchical network of smaller actions. OpenAI’s Operator (OpenAI, 2025b) allows users to “take control” of the agent to perform tasks manually. Yet, interactive techniques in this area are still nascent, despite the new challenges identified for human-agent communication (Bansal et al., 2024). Interfaces for LLM assistants (e.g., AgentGPT (Reworkd, 2024), Devin (Wu, 2024)) have been primarily limited to chat interfaces targeted at monitoring agent activity rather than empowering the user to proactively collaborate with the agent. An emerging body of work has started to experiment with more interactive techniques for agents. Kazemitabaar et al. (Kazemitabaar et al., 2024) developed interfaces for data analysts to edit an execution plan of an LLM to provide more control points for steering behavior, while Google’s Deep Research (Google, 2025) can re-generate a plan if the user provides feedback via prompting. However, these interactions are corrective rather than collaborative—that is, users would typically engage in these interactions to correct agent behavior post-hoc rather than proactively iterate back-and-forth with the agent at multiple points in the workflow. We target the latter in our work to elevate human feedback to beyond simply a corrective mechanism. That is, our goal is to enable interactive agents that effectively elicit and incorporate key guidance from the user through co-created human-AI representations.
@@ 2.2. Computational Notebooks
The computational notebook is an interactive paradigm that organizes program imperatives, input data, and rich outputs (e.g., tables, visualizations, interactive widgets) into linearly arranged cells (Chattopadhyay et al., 2020; Lau et al., 2020), reflecting Donald Knuth’s early visions of literate programming (Knuth, 1984). Computational notebooks are a well-studied paradigm in HCI (e.g., (Choi et al., 2023; Ayobi et al., 2023; Zheng et al., 2022; Lin et al., 2023; Chattopadhyay et al., 2020; O’Leary et al., 2023)), with most prior work focusing on how data scientists leverage them in their workflows.
In our work, we take inspiration from design decisions made for computational notebooks because of the many parallels between workflows in notebooks and mixed-initiative task completion systems (Horvitz, 1999; Grunde-McLaughlin et al., 2023; Wu et al., 2023). For instance, data scientists often break down a high-level data analysis task into executable cells in computational notebooks (Kery et al., 2018; Rule et al., 2018). As they execute each cell and inspect outputs, they can verify the quality and sensibility of outputs before moving forward in their analysis, or go back to modify existing code cells before execution to interleave between programming, execution, and output examination (Kery et al., 2020).
@@
By drawing these parallels, we can reveal new opportunities to design interactive workflows and interfaces for LLM agents. Just like how a data scientist may easily add new cells in a notebook or reconfigure existing cells to adapt their analysis plan, a user may interactively edit an agent’s plan of execution to better steer the agent in productive directions. Furthermore, an agentic system may simultaneously be more usable and resource-efficient if a user could direct an agent to iterate on a subtask to improve its output before continuing onto subtasks dependent on that output. This also demands new ways of viewing and editing agent outputs on subtasks, which increases opportunities for human input over existing approaches such as simply logging agent actions (AutoGPT, 2024; Reworkd, 2024; OpenAI Platform, 2024). Our work exploits these parallels to contribute new design patterns for steering agentic systems.
2.3. AI for Scientific Research
Significant efforts in recent years have advanced how AI can be used to support scientific research. Of particular interest to us are works that leverage these recent advances to help researchers with planning and execution of literature-focused tasks. Planning a research project is a cognitively demanding, complex process consisting of iterative cycles of divergent and convergent thinking grounded in literature review (Palani et al., 2023; Dean et al., 2006; Chang et al., 2023), making it an ideal scenario to explore human-AI collaboration.
One major thread of prior research has focused on leveraging AI models as tools or components in user-driven systems. These include tools for paper reading (Chang et al., 2023; Rachatasumrit et al., 2022) and skimming (Fok et al., 2022), literature review (Lee et al., 2024b; Palani et al., 2023), paper recommendation (Kang et al., 2022b, a), information retrieval and sensemaking (Chan et al., 2018; Kang et al., 2023; Fok et al., 2023; Wang et al., 2024a), ideation (Gu and Krenn, 2024; Baek et al., 2024; Jansen et al., 2024) and “deep research” systems that combine support for multiple activities (Google, 2025; OpenAI, 2025a; Perplexity, 2025). Progress in this area has given rise to excitement for a “computational inflection for scientific discovery” (Hope et al., 2022). On the other end of the spectrum are fully automated end-to-end systems where an AI agent has much more agency in planning and executing tasks to attempt to carry out entire research projects on their own (Lu et al., 2024; Jansen et al., 2025; Gottweis et al., 2025).